OpenStudy (anonymous):

Andrew can paint the neighbor's house 2 times as fast as bailey. The year Andrew and bailey worked together, it took 1 days. How long would it take each to paint the house? I need help on this Please as soon as possible?

OpenStudy (anonymous):

ok lets use d = rt equation

OpenStudy (anonymous):

Andrew did twice the work Bailey did. So bailey painted 1/3 house and Andrew painted 2/3 in one day. So Bailey will take \[1\div1/3= 3 days\] Andrew will take \[1\div2/3=3/2 days = 1.5 days\]

OpenStudy (anonymous):

hope that halped